OpenStudy (anonymous):

A cone has a base radius of 5 centimeters and a height of 18 centimeters. What is the volume of the cone? A.450π cm3 B.150π cm3 C.90π cm3 D.30π cm3

OpenStudy (anonymous):

Formula: \(V = \dfrac{1}{3} \pi r^2 h\) Plug in what we know: \(V = \dfrac{1}{3}\pi (5)^2 (18)\) What's 5 * 5?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Yes, and what's 25 * 18?

OpenStudy (anonymous):

450

OpenStudy (anonymous):

Yes, and what's 1/3 * 450?

OpenStudy (anonymous):

ohhh 150